Calculate the mass of an object if it moves down an inclined plane of 45 degrees with a force of 70N. Treat Acceleration due to gravity as 6.0ms and ignore the effects of friction in this calculation.Β 


1
Expert's answer
2021-05-16T17:59:08-0400

"mg\sin45°-70=ma\to m=\frac{70}{g\sin45°-a}=\frac{70}{9.8\cdot\sin45°-6}=75.3\ (kg)"